OpenStudy (anonymous):

Write an equation of a line in slope intercept form that is parallel to y=3x+6 and passes through the point (-10, 2.5)

OpenStudy (anonymous):

Let us first find the slope of y=3x+6 comparing y=3x+6 with y=mx+c we find slope m= 3 and y-intercept c = 6 Since parallel lines have same slope. therefore slope (m) of the required line is also = 3 since line passes through the point (-10, 2.5) Equation of line in slope -point form is given as; \[\huge y-y_1 =m(x-x_1)\] where \[\huge x_1 =-10\] & \[\huge y_1 =2.5\] Now substituting the values of m, x1 & y1 in the above equation we find; \[\huge y-2.5 =3[x-(-10)]\] \[\huge y-2.5 =3[x+10] \rightarrow \huge y-2.5 =3x+30\] \[\rightarrow \huge y =3x+30+2.5 \rightarrow \huge y =3x+32.5\] Hence y =3x+32.5 is the required equation of the line. @sky6298

OpenStudy (anonymous):

@sky6298

OpenStudy (anonymous):

Thank you!